Aqueous sulfuric acid H

2

SO

4

 reacts with solid sodium hydroxide NaOH

 to produce aqueous sodium sulfate Na

2

SO

4

 and liquid water H

2

O

. If 9.14

g

 of water is produced from the reaction of 85.3

g

 of sulfuric acid and 36.9

g

 of sodium hydroxide, calculate the percent yield of water.


1
Expert's answer
2021-11-05T01:07:27-0400

Write a balanced chemical equation;

"H_2SO_4+2NaOH \to Na_2SO_4+2H_2O"

moles of "H_2SO_4" "=85.3\/98"

moles of "NaOH=36.9\/40"

"NaOH" has less moles hence it is the limiting reagent.

Now 2moles of "NaOH" = 2moles of "H_2O" formed.

Hence moles of water formed "=36.9\/40"

Mass of water produced (theoretical) "=\frac{36.9}{40}\times 18=16.605g"

Actual mass of water produced "=9.14g"

"\%yield=\frac{9.14}{16.605}\times 100\%=55.04\%"